Find three consecutive integers where 4 more then twice the first is 12 more than the third.
Answer by stanbon(48558)

Find three consecutive integers where 4 more then twice the first is 12 more than the third.
----------------
1st: x
2nd: x+1
3rd: x+2
EQUATION:
2x+4=x+2+12
x = 10
x+1 = 11
x+2 = 12
===========
Cheers,
Stan H.